# EnvSDD: Benchmarking Environmental Sound Deepfake Detection

### v1

@@ 1 Introduction
Several studies have focused on fake environmental sound detection. In SceneFake [7], the authors first enhance the real speech involving a scene (e.g., “Airport”), then add another scene to the enhanced speech (e.g., “Street”), resulting in a scene-fake audio clip. Similar to SceneFake, in EnvFake [8], the authors create a multimodal scene-fake dataset by combining visual and audio information. These works focus on scene-consistency detection, where all environmental audio clips are sourced from real-life scenarios rather than being generated by AI models. Ouajdi et al. [9] use real-fake paired audio data from a Foley sound synthesis challenge for ESDD, where fake sounds are generated by 44 submitted audio generation systems. However, this work only includes monophonic audio, and only conducts in-domain evaluation, without considering the models’ generalization to unseen domains. In FakeSound [10], the authors mask a portion of the audio and use audio inpainting models to generate new audio in the masked segments, creating fake sound clips. In this work, most audio clips are polyphonic, with limited consideration of monophonic conditions. In addition, they use a relatively small subset of AudioCaps [11] for generating fake audio, with just about 5 hours of data.
@@ 2.1 Real Data Collection
Monophonic datasets: UrbanSound8K is a widely used dataset for environmental sound classification, specifically designed for urban sound recognition tasks, which contains 10 different urban sound classes, such as “siren” and “drilling”. We select audio clips with a duration of 4 seconds and a sampling rate above 16 kHz from it, resulting in 4,523 clips. DCASE 2023 Task7 Dev refers to the development set of Task 7 from DCASE 2023 challenge, which contains real-life audio clips from various public resources with 7 event classes.
Polyphonic datasets: TAU UAS 2019 Open Dev is designed for acoustic scene classification, consisting of audio clips recorded in 14 different real-life acoustic scenes, such as “airport” and “office”. TUT SED 2016 and TUT SED 2017 are two environmental audio datasets for sound event detection, containing real-life recordings from various indoor and outdoor scenarios. Clotho is a dataset designed for audio captioning tasks, which contains audio clips from FreesSound [20], and human-written captions describing the audio content.
Since DCASE 2023 Task 7 Dev and Clotho are used only for evaluation, we select a small subset from these datasets, covering various real-life scenarios. Additionally, the “DogBark” and “GunShot” events in DCASE 2023 Task 7 Dev overlap with UrbanSound8K, and are therefore excluded.
@@ 2.2 Deepfake Data Generation
As shown in Figure 1, for monophonic audio, we directly rewrite the metadata as the audio caption. “Rewrite” refers to removing meaningless characters from the metadata, capitalizing the first character, and adding a period in the end (e.g., “gun_shot” is rewritten to “Gun shot.”). For polyphonic audio, a large language model (LLM) is applied to generate captions based on the metadata. Specifically, we use Mistral 7B [24] as the LLM, with prompts configured as follows. If the metadata contains both scene and event labels, the prompt is “Prompt A” here as shown in the box:
@@ 4.1 Experimental Setup
Table 4: Performance in EER (%) of baseline and proposed systems on EnvSDD dataset. “Seen SD” means “Seen Source Datasets” and “Seen GM” means “Seen Generation Model”.
@@ 4.2 Performance of Baselines and Proposed System
Let us consider having a closer look at the performances of all the systems under each test condition. The low EERs (below 1%) for all systems in Test 01 indicates that in-domain sound deepfake detection is a relatively simple task. In Test 02, the trained baselines have to handle fake data generated by unseen TTA and ATA models. The results show a significant drop in performance on Test 02 compared to Test 01, demonstrating the detection model’s limited ability to generalize across different audio generation models. Similarly, in Test 03, the source datasets are out-of-domain for the trained detection models, leading to degraded performance compared to Test 01. Furthermore, we observe that AASIST and BEATs+AASIST systems are more affected by unseen source datasets than unseen audio generation models, whereas W2V2+AASIST shows an opposite trend indicating more dependence on generation models. Test 04 is the most challenging condition as both the audio generation models and the source datasets are out-of-domain, resulting significant drop in performance for all the systems. However, the proposed BEATs+AASIST system still performs substantially better than the two baselines.

### v2

@@ 1 Introduction
Several studies have focused on fake environmental sound detection. In SceneFake [7], the authors first enhance the real speech involving a scene (e.g., ``Airport''), then add another scene to the enhanced speech (e.g., ``Street''), resulting in a scene-fake audio clip. Similar to SceneFake, in EnvFake [8], the authors create a multimodal scene-fake dataset by combining visual and audio information. These works focus on scene-consistency detection, where all environmental audio clips are sourced from real-life scenarios rather than being generated by AI models. Ouajdi et al. [9] use real-fake paired audio data from a Foley sound synthesis challenge for ESDD, where fake sounds are generated by 44 submitted audio generation systems. However, this work only includes monophonic audio, and only conducts in-domain evaluation, without considering the models’ generalization to unseen domains. In FakeSound [10], the authors mask a portion of the audio and use audio inpainting models to generate new audio in the masked segments, creating fake sound clips. In this work, most audio clips are polyphonic, with limited consideration of monophonic conditions. In addition, they use a relatively small subset of AudioCaps [11] for generating fake audio, with just about 5 hours of data.
@@ 2.1 Real Data Collection
Monophonic datasets: UrbanSound8K is a widely used dataset for environmental sound classification, specifically designed for urban sound recognition tasks, which contains 10 different urban sound classes, such as ``siren'' and ``drilling''. We select audio clips with a duration of 4 seconds and a sampling rate above 16 kHz from it, resulting in 4,523 clips. DCASE 2023 Task7 Dev refers to the development set of Task 7 from DCASE 2023 challenge, which contains real-life audio clips from various public resources with 7 event classes.
Polyphonic datasets: TAU UAS 2019 Open Dev is designed for acoustic scene classification, consisting of audio clips recorded in 14 different real-life acoustic scenes, such as ``airport'' and ``office''. TUT SED 2016 and TUT SED 2017 are two environmental audio datasets for sound event detection, containing real-life recordings from various indoor and outdoor scenarios. Clotho is a dataset designed for audio captioning tasks, which contains audio clips from FreesSound [20], and human-written captions describing the audio content.
Since DCASE 2023 Task 7 Dev and Clotho are used only for evaluation, we select a small subset from these datasets, covering various real-life scenarios. Additionally, the ``DogBark'' and ``GunShot'' events in DCASE 2023 Task 7 Dev overlap with UrbanSound8K, and are therefore excluded.
@@ 2.2 Deepfake Data Generation
As shown in Figure 1, for monophonic audio, we directly rewrite the metadata as the audio caption. ``Rewrite'' refers to removing meaningless characters from the metadata, capitalizing the first character, and adding a period in the end (e.g., ``gun_shot'' is rewritten to ``Gun shot.''). For polyphonic audio, a large language model (LLM) is applied to generate captions based on the metadata. Specifically, we use Mistral 7B [24] as the LLM, with prompts configured as follows. If the metadata contains both scene and event labels, the prompt is ``Prompt A'' here as shown in the box:
@@ 4.1 Experimental Setup
Table 4: Performance in EER (%) of baseline and proposed systems on EnvSDD dataset. ``Seen SD'' means ``Seen Source Datasets'' and ``Seen GM'' means ``Seen Generation Model''.
@@ 4.2 Performance of Baselines and Proposed System
Let us consider having a closer look at the performances of all the systems under each test condition. The low EERs (below 1%) for all systems in Test 01 indicates that in-domain sound deepfake detection is a relatively simple task. In Test 02, the trained baselines have to handle fake data generated by unseen TTA and ATA models. The results show a significant drop in performance on Test 02 compared to Test 01, demonstrating the detection model's limited ability to generalize across different audio generation models. Similarly, in Test 03, the source datasets are out-of-domain for the trained detection models, leading to degraded performance compared to Test 01. Furthermore, we observe that AASIST and BEATs+AASIST systems are more affected by unseen source datasets than unseen audio generation models, whereas W2V2+AASIST shows an opposite trend indicating more dependence on generation models. Test 04 is the most challenging condition as both the audio generation models and the source datasets are out-of-domain, resulting significant drop in performance for all the systems. However, the proposed BEATs+AASIST system still performs substantially better than the two baselines.